# AdTEC: A Unified Benchmark for Evaluating Text Quality in Search Engine Advertising

### v1

@@ Abstract
With the increase in the more fluent ad texts automatically created by natural language generation technology, it is in the high demand to verify the quality of these creatives in a real-world setting. We propose AdTEC, the first public benchmark to evaluate ad texts in multiple aspects from the perspective of practical advertising operations. Our contributions are: (i) Defining five tasks for evaluating the quality of ad texts and building a dataset based on the actual operational experience of advertising agencies, which is typically kept in-house. (ii) Validating the performance of existing pre-trained language models (PLMs) and human evaluators on the dataset. (iii) Analyzing the characteristics and providing challenges of the benchmark. The results show that while PLMs have already reached the practical usage level in several tasks, human still outperforms in certain domains, implying that there is significant room for improvement in such area.
@@ 1 Introduction
Online advertising, especially sponsored search advertising, is a dominant sector for vendors to promote their products, and the market size is estimated to grow by billions of dollars over the next few years [1]. To meet the increasing demands of advertising operations (AdOps), such as creating ad texts from product information (Step 2 in Figure 1), the remarkable success of natural language generation (NLG) by pre-trained language models (PLMs) [2, 3, 4, 5] has given a boost to applications in practice [6, 7, 8], making advertising a huge industrial use case for NLP.
In this paper, we focus on evaluating the quality of ad texts generated by such models. We refer to this process as ad text evaluation, as depicted in Step 3 of Figure 1. Ad text evaluation is crucial because low-quality ad texts can disadvantage advertisers through a lack of fluency, inappropriate appeals, and misleading representations. Because it is expensive and non-scalable to ask humans for verifying the quality of each text in high-volume domains such as sponsored search advertising, it is in the high demand for developing automatic quality estimators for ad texts. The quality has multiple dimensions, such as appropriate wording, effective appeals, consistency between ad text and product information, and high predicted performance. Although these dimensions should be included for evaluating automatic quality estimators, there is no such benchmark. Consequently, the bottleneck lies in verifying the quality of ad texts despite the ability of generating numerous creatives automatically, which hinders the scalability of delivery volume. Thus, we aim to construct a benchmark for evaluating the quality of ad texts.
The primary challenge in constructing an ad text evaluation benchmark is the absence of a clear definition for the tasks [9, 10]. The lack of domain knowledge in AdOps complicates the understanding of high-quality ad text standards and the accurate definition of tasks. However, the AdOps workflow is complex, relies on various platforms and formats, and encompasses multiple methodologies and metrics. In addition, only a few companies possess the expertise to operate online advertising at scale. Owing to legal and contractual obligations, advertising workflows and data are predominantly managed in-house, leading to a lack of publicly available datasets. This scarcity makes it challenging to systematically reproduce and validate diverse methodologies in academia. Consequently, research remains less active in the advertising field, potentially overlooking issues and delaying the application and development of cutting-edge technology.
In response to these challenges, we propose AdTEC (Ad Text Evaluation Benchmark by CyberAgent), the first publicly available benchmark that defines and unifies tasks based on generalized AdOps workflows. Our major contributions are as follows:
@@
We carefully designed five tasks that closely match real-world advertising workflows based on practical operational experiences of advertising agencies. Based on these tasks, we constructed a high-quality Japanese dataset from the operational data and released it publicly, even though such data is typically being kept in-house and difficult to obtain. These datasets are available at https://github.com/CyberAgentAILab/AdTEC.
@@
We analyzed the characteristics and identified potential issues of the dataset through experiments, demonstrating that our benchmark is challenging, and highlighting potential areas for improvement and future research.
2 Understanding Sponsored Search Advertising and its Workflow
Sponsored search advertising displays titles and descriptions of ads that are relevant to keywords, which users enter into search engines, appearing as part of the search results. As illustrates in Figure 2, when a user clicks ad’s URL, they are directed to a web page known as a landing page (LP). LP contains texts and images related to the advertiser’s products, allowing the user to take actions, such as making a purchase.
Operating such ads requires high expertise, owing to the variety of delivery platforms, properties and formats. To gain insights, we interviewed two types of experts familiar with AdOps; those overseeing AdOps departments at ad agencies and those directly involved in on-site operations. Through these interviews, we explored the intricacies of AdOps, generalized the workflow, and delineated six steps, as illustrated in Figure 1. In this work, we focus mainly on Step 3 of quality evaluation.
3 Building AdTEC
Our goal is to develop a benchmark that captures multi-dimensional aspects of quality that are relevant to practical scenarios and carefully curated to provide value for both real-world applications and various research purposes.
3.1 Task Design
We followed the workflow outlined in §2, and asked the same experts from the interview to identify crucial parts within Step 3 in Figure 1. Derived from their insights, we generalized and defined tasks that adhere to the principle of evaluating ad texts directly or indirectly. Direct evaluation tasks are used to test texts with strict criteria, such as a binary pass/fail outcome or a numerical score to quantify the text quality. These tasks will serve as a checklist to ensure that minimum delivery standards are met. Indirect evaluation tasks are used to assist human evaluators in reviewing or refining texts or serve as a bridge to connect the downstream tasks. Based on the above principles, we designed five tasks: three direct evaluation tasks (Ad Acceptability, Ad Consistency, and Ad Performance Estimation) and two indirect evaluation tasks (A3 Recognition and Ad Similarity).111Note that all examples in this paper are translated from Japanese into English for the presentation brevity.
@@ 3.1 Task Design
Ad Acceptability . As most ad delivery platforms impose text length restrictions, minor grammatical errors are tolerated to enhance the readability and engage readers within limited space. However, excessive compression can mislead readers, and such poor-quality ads should be detected before delivery to avoid negative impacts on the advertiser. To assess this, we defined the task of Ad Acceptability, which predicts the acceptance of overall quality with binary labels: acceptable/unacceptable. Unacceptable ad phenomena include “collapsing symbols,” “unnatural repetition,” and “incomprehensible meaning” based on expert feedback. This differs from the general concept of linguistic acceptability, which checks for grammatical correctness, such as CoLA [11]. Examples of Ad Acceptability are shown in Table 1. The ad text “Engineer’s career / Engineer’s Jobs” is unacceptable because the meaning is duplicated.
Ad Consistency . Verifying consistency between the ad text and LP content is crucial. If a feature or price mentioned in the ad text is not referenced in the corresponding LP, it may violate the Law for the Prevention of Unjustified Extra or Unexpected Benefit and Misleading Representation, resulting in damages to the advertiser. However, these inconsistencies are hard to detect as some factual expressions do not appear in LPs. For example, the term “official,” which is often used in ad text but rarely appears in LP content. To assess this, we defined the Ad Consistency task, which predicts the consistency between LP content and ad text with binary labels: consistent/inconsistent. Examples of Ad Consistency are shown in Table 1. The first line is labeled inconsistent because the LP refers to a “single-family home,” while the ad text mentions a “condominium.”
Ad Performance Estimation . The most straightforward way to measure ad quality is to publish them online and let end customers evaluate them. However, delivering all ads without alterations is impractical, as low-quality ads can negatively impact advertisers. Therefore, prior studies investigated offline methods to measure ad text quality by simulating customer behavior, such as click-through rate (CTR), based on past delivery history. These methods are currently standard practices in many organizations. Inspired by their works [12, 13, 14], we adopted Ad Performance Estimation task to estimate a quality score in the range of [0, 100] from ad texts, keywords, and industry, as shown in Table 2. The score simulates customer behavior based on past delivery history, and is non-linearly transformed to maintain the original label distribution for contractual reasons.
A3 Recognition . One of the most crucial factors in advertising is the aspect of advertising appeals (A3). At its core, advertising aims to connect advertisers with readers, and A3 serves as a bridge between them. For example, an ad emphasizing low cost may resonate with price-conscious readers, while one focusing on high performance may not. Thus, recognizing appealing expressions in advertising and using appropriate A3 can enhance downstream tasks, such as CTR prediction [9]. In the A3 Recognition task, we follow a previous study [9], which predicts all relevant A3 labels in a given ad text. Figure 3 shows an example of ad texts and the corresponding A3s. All labels and distributions can be found in Appendix Table 13.
Ad Similarity . Repeatedly showing same ads to readers leads to ad fatigue [15], where readers become bored and ad performance declines. Therefore, it is essential to avoid displaying the same ads for extended periods, and regularly replace them with different ones. However, the transition from old to new ads must be carefully managed because we need to maintain the product and its appeal while updating the wording or representations, or we risk disengaging customers who were attracted to the previous ads. Thus, measuring the similarity particularly focusing on this situation is crucial, which enables us to determine whether to replace the ad based on a quantified score.
Building on the aforementioned motivation, we defined the Ad Similarity task, which predicts the similarity score for an ad text pair on a scale of [1, 5]. The lower the values, the less similar the pair, and vice versa. Examples are shown in Table 3. The first example pair illustrates high similarity, where the both the product of “Suppon Black Vinegar” and the A3 of luxury are identical. Conversely, the second example pair differ in A3 with budget and discount, resulting in relatively low similarity.
@@
| Field Example value Title 1 [No.1] Card loan comparison site Title 2 A must-see for those who in a hurry! Title 3 Instant Loan Secure Card Loan Desc. 1 The best place to get a card loan without telling anyone. You only need a driver’s license to apply Desc. 2 Convenient to use ATMs at convenience stores. Convenient and quick loans are available if you apply before 10:00 p.m. Keyword card loan Industry finance Score 82.3 Table 2: Examples of Ad Performance Estimation task. Desc. represent Description. Score is the label and others are inputs. |
@@ 3.2 Dataset Construction
Data Collection. For the Ad Acceptability and Ad Consistency task, data was collected from the ad creation phase of the actual AdOps workflow, including both human creators and NLG models’ outputs. For the Ad Performance Estimation and Ad Similarity task, we used Japanese sponsored search ads delivered between 2021 and 2022. For the A3 Recognition task, we used data from Murakami et al. [9].
Data Pre-processing. In the Ad Performance Estimation task, we negotiated with our clients to use a subset of the data approved for public release, as the data includes ad performance metrics sensitive to advertisers. Furthermore, we applied a nonlinear transformation to the raw CTR, scaling them to the range of to preserve their distribution. Additionally, we masked proper nouns (e.g., product names, company names) to prevent identification of advertisers and avoid any potential negative repercussions upon data release. In the Ad Similarity task, creating sentence pairs through random sampling is inefficient because most pairs are not similar. Therefore, we utilized the account structure configured during ad delivery, which includes client, account, campaign, ad group, and keyword information, as depicted in Figure 4. We created pseudo-similar pairs by sampling texts from the same ad group, as preliminary results suggested. To balance the label distribution, we created pseudo-dissimilar pairs by ensuring that the two texts belonged to different clients. Consequently, we sampled pseudo-similar and pseudo-dissimilar pairs at a ratio of 9:1.
Annotation Workflow. All annotators are native Japanese speakers and experts with vast experience in AdOps. The annotation workflow for all tasks, except Ad Performance Estimation and A3 Recognition, is as follows: (1) We first removed duplicate entries and filtered out data in languages other than Japanese. (2) We iteratively revised the annotation guidelines until we achieved a satisfactory level of agreement through pilot annotations on small sampled datasets. (3) Following the guidelines, we conducted a pilot annotation by asking three annotators to annotate the same sampled data used in Step 2. Thereafter, we compared the annotators’ results with our own and resolved any inconsistencies to further refine the guidelines. This cycle was repeated at least twice. (4) After completing Steps 1 to 3, we conducted the main annotation on the full test set using the finalized guidelines. The complete guidelines are provided in the Appendix.
@@
Data Splitting. Despite efforts at deduplication, similar ad expressions can still be easily found, potentially leading to data leakage with a simple random split. Furthermore, assuming the data is used in industry, it is crucial to generalize effectively without overfitting to specific ad expressions. Therefore, for Ad Acceptability and Ad Consistency, we split the data, considering the ad hierarchical structure in Figure 4, ensuring that clients do not overlap across training, development, and testing. For the Ad Performance Estimation, we used the delivery structure, with the non-overlapping layer as the campaign. For the A3 Recognition, we used the same split from Murakami et al. [9]. For the Ad Similarity, we randomly split the data to maintain label distribution consistency.
4 Benchmark Experiment Settings
In this section, we conducted an experiment to explore the performance on the proposed benchmark.
4.1 Evaluation Metrics
Table 5 provides a brief overview of each task and the corresponding metrics used to measure task performance. Our tasks are categorized into three setups: binary classification, multi-label classification, and regression. For binary classification, we use accuracy and the F1-score to evaluate the binary labels. In multi-label classification, we follow Murakami et al. [9] and use the F1-score with both macro and micro settings. For regression, we use Pearson and Spearman correlation coefficients. These metrics are standard practices in numerous studies.
| Task | Setup: Input Label | Metrics |
| Ad Similarity | Regression: Ad text pair | Pearson/Spearman corr. |
4.2 Evaluators
We employed two types of evaluators: the PLMs, including both fine-tuned settings with encoder models and zero-/few-shot settings for LLMs, as well as human evaluators. Please note that fine-tuning an LLM is beyond the scope, owing to the cost constrains.
Fine-tuning Setting with Encoder Models.
In the fine-tuning settings, models receive ad text and other values as input and predict labels , where represents the -th token of the input sequence. The hyperparameters used are shown in Appendix Table 12. Ad Acceptability and Ad Consistency involve predicting binary label using an MLP from the encoded vector representation , where is the special token for classification. The inputs of the Ad Acceptability task only use the ad text , while Ad Consistency uses as input, where and represent the LP text and concatenation by a special token , respectively. A3 Recognition involves predicting all possible labels from a given ad text, . We adopted the doc-based architecture described in Murakami et al. [9], where the encoder model was used to obtain the vector representation . This vector was then fed into an MLP layer, producing a label probability distribution , where and is 21. This number corresponds to the number of A3 labels defined by Murakami et al. [9], as shown in Appendix Table 13. Ad Similarity and Ad Performance Estimation are regression tasks that involve predicting a value of range and , respectively. Similar to the text classification task, we concatenated all inputs for each task with the special token , encoded the input into vector representation , and then fed the vector into the MLP layer to predict .
We utilized publicly available encoder models as baselines, namely Tohoku BERT, Waseda RoBERTa, and XLM-RoBERTa [16], commonly employed for Japanese NLP tasks, which differ in pre-tokenizer, tokenization unit, and pre-training dataset. All the aforementioned models are of LARGE size, with detailed information provided in Appendix Tables 8 and 9. We conducted experiments using the BASE size model, however, due to space constraints, the results from BASE size models are provided in Appendix Table 10.
Zero-/Few-shot Setting with LLMs.
We employed CALM27b and ELYZA7bas baselines for open LLMs. CALM27b and ELYZA7b are based on the Llama 2 [17] architecture but differ in training methods and data; CALM27b was trained from scratch, whereas ELYZA7b was continuously trained from the original Llama 2. These models were chosen because of their public availability and strong performance. Additionally, OpenAI’s GPT-3.5 and GPT-4 were included as reference baselines, owing to their exceptional performance, despite not being publicly available. Note that the results from GPT-3.5 are shown in the complete result in Appendix Table 10.
Before conducting the experiment, we sampled 100 instances from the development set to adjust the parameters and prompts. The prompts we used are shown in Appendix Figure 7. Additionally, we determined the number of shots that performed best in the development set: 3-shot for Ad Similarity and Ad Performance Estimation task, 2-shot for A3 Recognition, and zero-shot for the others. We calculated the final score by averaging five runs with different few-shot examples for a single instance.
Human.
We enlisted three human evaluators who are not engaging in AdOps to evaluate all tasks except Ad Performance Estimation. We followed the same procedure for instruction as described in §3.2. Pilot evaluations were carried out twice on randomly sampled 100 instances from the training set for each task before the main run. In the final assessment, a majority vote per instance was conducted for Ad Acceptability and Ad Consistency tasks, while the average scores of evaluator assessments were reported for other tasks.
@@ 4.2 Evaluators
| Human | 0.732/0.790 | 0.703/0.807 | — | 0.564/0.538 | 0.699/0.765 |
5 Result and Discussion

### v2

@@ Abstract
As the fluency of ad texts automatically generated by natural language generation technologies continues to improve, there is an increasing demand to assess the quality of these creatives in real-world setting. We propose AdTEC, the first public benchmark to evaluate ad texts from multiple perspectives within practical advertising operations. Our contributions are as follows: (i) Defining five tasks for evaluating the quality of ad texts, as well as constructing a Japanese dataset based on the practical operational experiences of advertising agencies, which are typically maintained in-house. (ii) Validating the per-formance of existing pre-trained language models (PLMs) and human evaluators on this dataset. (iii) Analyzing the characteristics and providing challenges of the benchmark. Our results show that while PLMs have a practical level of performance in several tasks, humans continue to outperform them in certain domains, indicating that there remains significant potential for further improvement in this area. The dataset is publicly available at: https://cyberagentailab.github.io/AdTEC.
AdTEC: A Unified Benchmark for Evaluating Text Quality
in Search Engine Advertising
Peinan Zhang, Yusuke Sakai, Masato Mita, Hiroki Ouchi, Taro Watanabe CyberAgent, Nara Institute of Science and Technology {zhang_peinan,mita_masato}@cyberagent.co.jp {sakai.yusuke.sr9,hiroki.ouchi,taro}@is.naist.jp
@@ 1 Introduction
Online advertising, especially sponsored search advertising (Figure 1), is a dominant sector for vendors to promote their products, and the market size is estimated to grow by billions of dollars over the next few years Murakami et al. (2023). To meet the increasing demands of advertising operations (AdOps), such as creating ad texts from product information (Step 2 in Figure 2), the remarkable success of natural language generation (NLG) by pre-trained language models (PLMs) Dong et al. (2021); Vaswani et al. (2017); Brown et al. (2020); Touvron et al. (2023a) has given a boost to practical applications Hughes et al. (2019); Kamigaito et al. (2021); Golobokov et al. (2022), making advertising a huge industrial use case for natural language processing (NLP).
In this paper, we focus on evaluating the quality of ad texts generated by such models. We refer to this process as ad text evaluation, as depicted in Step 3 of Figure 2. Ad text evaluation is crucial because low-quality ad texts , which may lack fluency, present inappropriate appeals, or convey misleading representations, can negatively impact advertisers. Since it is costly and not scalable to have humans verify the quality of each text in high-volume domains, such as sponsored search advertising, there is a high demand for the development of automated quality estimators for ad texts. The quality has multiple dimensions, such as appropriate wording, effective appeals, consistency between ad text and product information, and high predicted performance. Although these dimensions should be included when evaluating automatic quality estimators, no benchmark currently exists to assess them. Consequently, the bottleneck lies in verifying the quality of ad texts despite the ability to generate numerous creatives automatically, which hinders delivery volume scalability. Thus, we aim to construct a benchmark to evaluate the quality of ad texts.
The primary challenge in constructing an ad text evaluation benchmark is the absence of a clear definition for tasks Murakami et al. (2022); Mita et al. (2024). The lack of domain knowledge in AdOps complicates the understanding of high-quality ad text standards and the accurate definition of tasks. However, the AdOps workflow is complex, relies on various platforms and formats, and encompasses multiple methodologies and metrics. In addition, only a few companies possess the expertise to operate online advertising at scale. Owing to legal and contractual obligations, advertising workflows and data are predominantly managed in-house, leading to a lack of publicly available datasets. This scarcity makes it challenging to systematically reproduce and validate diverse methodologies in academia. Consequently, research activity is limited in the advertising field, which leaves potential issues unaddressed and delays the application and development of cutting-edge technology.
To address these challenges, we propose AdTEC, the first public benchmark that defines and unifies tasks based on generalized AdOps workflows. Our major contributions are as follows:
@@
We organized real-world advertising workflows and carefully designed five tasks to evaluate the quality of ad texts. This was based on the practical operational experiences of advertising agencies, using Japanese data as a case study. We are releasing this dataset111https://cyberagentailab.github.io/AdTEC, marking it as the first publicly available dataset for ad text evaluation, as such data are typically maintained in-house and difficult to obtain.
@@
We analyzed the characteristics and identified potential issues of the dataset through experiments, demonstrating that our benchmark is challenging and highlighting potential areas for improvement and future research.
@@
2 Understanding Sponsored Search Advertising and its Workflow
In sponsored search advertising, titles and descriptions of ads that are relevant to keywords entered into search engines by customers appear as part of the search results. As shown in Figure 1, when a customer clicks an ad’s URL, they are directed to a web page known as a landing page (LP). The LP contains texts and images related to the advertiser’s products, prompting the customer to view the product or make purchase decisions.
Operating such ads requires high expertise, owing to the variety of delivery platforms, properties, and formats. To gain insights, we interviewed two types of experts familiar with AdOps: those overseeing AdOps departments at ad agencies and those directly involved in on-site operations. Through these interviews, we explored the intricacies of AdOps, generalized the workflow, and delineated six steps, as illustrated in Figure 2. In this work, we focus mainly on Step 3 of quality evaluation.
3 Constructing AdTEC
Our goal is to develop a benchmark that captures multi-dimensional aspects of quality that are relevant to practical scenarios. In addition, these aspects will be carefully curated to provide value for both real-world applications and various research purposes.
3.1 Task Design
We followed the workflow outlined in §2 and asked the same experts from the interview to identify crucial parts within Step 3 in Figure 2. Based on their insights, we generalized and defined tasks that adhere to the principle of evaluating ad texts either directly or indirectly. Direct evaluation tasks are used to test texts against strict criteria, such as a binary pass/fail outcome or numerical score to quantify the text quality. These tasks serve as a checklist to ensure that minimum delivery standards are met. Indirect evaluation tasks are used to assist human evaluators in reviewing or refining texts or serve as a bridge to connect downstream tasks. Based on the above principles, we designed five tasks: three direct evaluation tasks (Ad Acceptability, Ad Consistency, and Ad Performance Estimation) and two indirect evaluation tasks (A3 Recognition and Ad Similarity).222Note that all examples in this paper are translated from Japanese into English for presentation brevity.
@@ 3.1 Task Design
| Field Example value Title 1 [No.1] Card loan comparison site Title 2 A must-see for those in a hurry! Title 3 Instant Loan Secure Card Loan Desc. 1 The best place to get a card loan without telling anyone. You only need a driver’s license to apply Desc. 2 Convenient to use ATMs at convenience stores. Convenient and quick loans are available if you apply before 10:00 p.m. Keyword card loan Industry finance Score 82.3 Table 2: Examples of Ad Performance Estimation task. Desc. represent Description. Score is the label, and others are inputs. |
Ad Acceptability.
As most ad delivery platforms impose text length restrictions, minor grammatical errors are tolerated to enhance the readability and engage customers within limited space. However, excessive compression can mislead customers, and such poor-quality ads should be detected before delivery to avoid negative impacts on the advertiser. To assess this, we defined the Ad Acceptability task, which predicts the acceptance of overall quality with binary labels: acceptable/unacceptable. Based on expert feedback, unacceptable ad phenomena include “collapsing symbols,” “unnatural repetition,” “incomprehensible meaning’, and “underspecified/vague”. This differs from the general concept of linguistic acceptability, which checks for grammatical correctness, such as CoLA Warstadt and Bowman (2019). Examples of Ad Acceptability are shown in Table 1. The ad text “Engineer’s career / Engineer’s Jobs” is unacceptable because the meaning is duplicated.
Ad Consistency.
Verifying consistency between the ad text and LP content is crucial. If a feature or price mentioned in the ad text is not referenced in the corresponding LP, it may violate the Law for the Prevention of Unjustified Extra or Unexpected Benefit and Misleading Representation, resulting in damages to the advertiser. However, these inconsistencies are difficult to detect as some factual expressions do not appear in LPs. For example, the term “official” is often used in ad text but rarely appears in LP content. To assess this, we defined the Ad Consistency task, which predicts the consistency between LP content and ad text with binary labels: consistent/inconsistent. Examples of Ad Consistency are shown in Table 1. The first line is labeled inconsistent because the LP refers to a “single-family home,” while the ad text mentions a “condominium.”
Ad Performance Estimation.
The most straightforward way to measure ad quality is to publish ads online and let end customers evaluate them. However, delivering all ads without alterations is impractical, as low-quality ads can negatively impact advertisers. Therefore, prior studies have investigated offline methods to measure ad text quality by simulating customer behavior, such as click-through rate (CTR), based on past delivery history. These methods are currently standard practices in many organizations. Inspired by such works Gharibshah et al. (2020); Niu and Hou (2020); Yang and Zhai (2022), we adopted the Ad Performance Estimation task to estimate a quality score in the range of [0, 100] based on ad texts, keywords, and industry, as shown in Table 2. The score simulates customer behavior based on past delivery history and is non-linearly transformed to maintain the original label distribution for contractual reasons.
A3 Recognition.
One of the most crucial factors in advertising is the aspect of advertising appeals (A3). At its core, advertising aims to connect advertisers with customers, and A3 serves as a bridge between them. For example, an ad emphasizing low cost may resonate with price-conscious customers, while one focusing on high performance may not. Thus, recognizing appealing expressions in advertising and using appropriate A3 can enhance downstream tasks, such as CTR prediction Murakami et al. (2022). In the A3 Recognition task, we follow a previous study Murakami et al. (2022), which predicted all relevant A3 labels in a given ad text. Figure 3 shows an example of ad texts and the corresponding A3s. All labels and distributions can be found in Table 11.
Ad Similarity.
Repeatedly showing the same ads to customers leads to ad fatigue Abrams and Vee (2007) and ad performance declines. Therefore, it is essential to avoid displaying the same ads for extended periods and regularly replace them with different ones. However, the transition from old to new ads must be carefully managed to maintain the product and its appeal. Specifically, while the wording and representations are being updated, there is a risk of disengagement of customers who were attracted to previous ads. Thus, when measuring similarity, we should particularly focus on this situation, which enables us to determine whether to replace the ad based on a quantified score.
Building on this motivation, we defined the Ad Similarity task, which predicts the similarity score for an ad text pair on a scale of [1, 5]. The lower the values, the less similar the pair, and vice versa. Examples are shown in Table 3. The first example pair illustrates high similarity, where both the product of “Suppon Black Vinegar” and the A3 of luxury are identical. Conversely, the second example pair differ in A3 with budget and discount, resulting in relatively low similarity.
@@ 3.2 Dataset Construction
| Task | Setup: Input Label | Metrics |
| Ad Similarity | Regression: Ad text pair | Pearson/Spearman corr. |
Data Collection.
For the Ad Accepta-bility and Ad Consistency tasks, data were collected from the ad creation phase of the actual AdOps workflow, including both human creators and NLG model outputs. For the Ad Performance Estimation and Ad Similarity tasks, we used Japanese sponsored search ads delivered between 2021 and 2022. For the A3 Recognition task, we used data from Murakami et al. (2022).
Data Pre-processing.
In the Ad Performance Estimation task, we negotiated with our clients to use a subset of the data approved for public release, as the data include ad performance metrics that are sensitive to advertisers. Furthermore, we applied a nonlinear transformation to the raw CTR, scaling the values to the range of to preserve their distribution. Additionally, we masked proper nouns (e.g., product and company names) to prevent identification of advertisers and avoid any potential negative repercussions upon data release. In the Ad Similarity task, creating sentence pairs through random sampling proved inefficient because most pairs are not similar. Therefore, we utilized the account structure configured during ad delivery, which includes client, account, campaign, ad group, and keyword information, as depicted in Figure 4. We created pseudo-similar pairs by sampling texts from the same ad group, as preliminary results suggested. To balance the label distribution, we created pseudo-dissimilar pairs by ensuring that the two texts belonged to different clients. Consequently, we sampled pseudo-similar and pseudo-dissimilar pairs at a ratio of 9:1.
@@
Annotation Workflow.
All annotators were native Japanese speakers and experts with at least two years of professional experience in AdOps. The annotation workflow for all tasks, except Ad Performance Estimation and A3 Recognition, followed these steps: (1) We first removed duplicate entries and filtered out data in languages other than Japanese. (2) We iteratively revised the annotation guidelines until we achieved a satisfactory level of agreement through pilot annotations on small sampled datasets. (3) Following the guidelines, we conducted a pilot annotation by asking three annotators to annotate the same sampled data used in Step 2. Thereafter, we compared the annotators’ results with our own and resolved any inconsistencies to further refine the guidelines. This cycle was repeated at least twice. (4) After completing Steps 1 to 3, we conducted the main annotation on the full test set using the finalized guidelines. The complete guidelines are provided in Appendix B.
Data Splitting.
@@
Despite efforts at deduplication, similar ad expressions can still be easily found, potentially leading to data leakage with a simple random split. Furthermore, assuming the data are used in industry, it is crucial to generalize effectively without overfitting to specific ad expressions. Therefore, for Ad Acceptability and Ad Consistency, we split the data, considering the ad hierarchical structure in Figure 4, ensuring that clients do not overlap across training, development, and testing. For Ad Performance Estimation, we used the delivery structure, with the non-overlapping layer as the campaign. For A3 Recognition, we used the same split from Murakami et al. (2022). For the Ad Similarity, we randomly split the data to maintain label distribution consistency.
@@
| Human | 0.732*/0.790* | 0.703/0.807* | — | 0.564/0.538 | 0.699/0.765 |
4 Experimental Settings
Table 4 provides a brief overview of each task and the corresponding metrics used to measure task performance. Our tasks are categorized into three setups: binary classification, multi-label classification, and regression. For binary classification, we use accuracy and F1-score to evaluate the binary labels. In multi-label classification, we follow Murakami et al. (2022) and use the F1-score with both macro and micro settings. For regression, we use Pearson and Spearman correlation coefficients.
We employed two types of evaluators: the PLMs, including both fine-tuned settings with encoder models and zero-/few-shot settings for LLMs, as well as human evaluators. Note that the detailed model description, hyperparameters, prompts, and additional information are provided in Appendix C.
Fine-tuning Setting with Encoder Models.
We utilized publicly available encoder models as baselines, specifically Tohoku BERT, Waseda RoBERTa, and XLM-RoBERTa Conneau et al. (2019), which are commonly employed for Japanese NLP tasks. These models differ in pre-tokenizer, tokenization unit, and pre-training dataset. All the aforementioned models are of LARGE size.
Zero-/Few-shot Setting with LLMs.
We employed CALM27b and ELYZA7bas baselines for open LLMs. CALM27b and ELYZA7b are based on the Llama 2 Touvron et al. (2023b) architecture but differ in training methods and data; CALM27b was trained from scratch, whereas ELYZA7b was continuously trained from the original Llama 2.333To further assess the gap between fine-tuning and zero-/few-shot learning of LLMs, we also fine-tuned an LLM. Due to the additional nature of this experiment, which falls beyond the scope of our study and is constrained by our computational resources, we selected only the best-performing open-source model from the zero-/few-shot settings. The settings and results are provided in Appendix C.
Human Evaluators.
We enlisted three human evaluators who are not involved in AdOps to evaluate all tasks except Ad Performance Estimation. We followed the same procedure for instruction as described in §3.2. Pilot evaluations were carried out twice on 100 randomly sampled instances from the training set for each task before the main run. In the final assessment, a majority vote per instance was conducted for the Ad Acceptability and Ad Consistency tasks, while the average scores of evaluator assessments were reported for other tasks.
5 Results and Discussion